# Entanglement witnesses and separability criteria based on generalized equiangular tight frames

### v3

@@ 5 Conclusions
In this paper, we applied the recently introduced generalized equiangular measurements to quantum entanglement detection. It was shown that, as long as GEAMs are conical 2-designs, the purity of the state is directly related to the partial index of coincidence, where the summation of squared probabilities is performed over few generalized equiangular tight frames. Such measurements were then used to define a class of positive but not trace-preserving maps and the corresponding entanglement witnesses. Interestingly, these witnesses are parametrized solely by two partial indices of coincidence. Next, we used the generalized equiangular measurements to construct separability criteria for bipartite states with subsystems of arbitrary dimensions.
In further research, it would be interesting to examine the properties of entanglement witnesses, such as extremality or optimality. Construction of witnesses and separability criteria from measurements that are not conical 2-designs remains as an open question.

### v4

@@ 6 Conclusions
In this paper, we applied the recently introduced generalized equiangular measurements to quantum entanglement detection. It was shown that, as long as GEAMs are conical 2-designs, the purity of the state is directly related to the partial index of coincidence, where the summation of squared probabilities is performed over few generalized equiangular tight frames. Such measurements were then used to define a class of positive but not trace-preserving maps and the corresponding entanglement witnesses. Interestingly, these witnesses are parametrized solely by two partial indices of coincidence. Examples for bound entanglement detection had been provided. Next, we used the generalized equiangular measurements to construct separability criteria for bipartite states with subsystems of arbitrary dimensions.
In further research, it would be interesting to examine the properties of entanglement witnesses, such as extremality or optimality. Construction of witnesses and separability criteria from measurements that are not conical 2-designs remains as an open question. Another possible direction is to generalize the presented formalism to multipartite systems and subsystems of unequal dimensionality.
@@ 7 Acknowledgements
8 Author Contribution
K.S. is solely responsible for the entire research process.